# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v3

@@ IV Conclusions
In this study, we conducted a detailed examination of buoyancy-aided mixed convective flow past a square cylinder at an angle of incidence, $\alpha=45^{\circ}$ for Re $=100$, $0.0\leq Ri\leq 1.0$ and Pr $=0.7$ by performing direct numerical simulations (DNS) using the finite-difference based Marker and Cell (MAC) method.
At a critical Ri between 0.6 and 0.7, vortex shedding suppression takes place, leading to a steady flow in the near-field. Simultaneously, the far-field develops large-scale unsteadiness similar to a plume. There is no range of Ri for which the entire flow field is seen to be steady. At low Ri, the downstream flow behavior resembles that of a wake, having a momentum deficit in the time-averaged velocity. As Ri is increased, the flow behavior gradually transforms into a jet-like plume, having a momentum excess in the time-averaged velocity. The flow retains the characteristics of both a wake and a plume in this intermediate Ri range. An inversion of vorticity takes place behind the cylinder for all positive Ri, because of the interplay between the forced and natural convective components of the flow. These findings provide valuable insight into the near-field and far-field dynamics of the buoyancy-aided mixed convective flows past cylinders.

### v4

@@ IV Conclusions
In this study, we conducted a detailed examination of buoyancy-aided mixed convective flow past a square cylinder at an angle of incidence, $\alpha=45^{\circ}$ for Re =100=100, $0.0\leq Ri\leq 1.0$ and Pr =0.7=0.7 by performing direct numerical simulations (DNS) using the finite-difference based Marker and Cell (MAC) method.
At a critical Ri between 0.6 and 0.7, vortex shedding suppression takes place, leading to a steady flow in the near-field. Simultaneously, the far-field develops large-scale unsteadiness similar to a plume. There is no range of Ri for which the entire flow field is seen to be steady. At low Ri, the downstream flow behavior resembles that of a wake, having a momentum deficit in the time-averaged velocity. As Ri is increased, the flow behavior gradually transforms into a jet-like plume, having a momentum excess in the time-averaged, which reveals a self-similar behavior. The flow retains the characteristics of both a wake (forced convective flow) and a plume (free convective flow) in this intermediate Ri range. An inversion of vorticity takes place behind the cylinder for all positive Ri, because of the interplay between the forced and natural convective components of the flow. These findings provide valuable insight into the near-field and far-field dynamics of the buoyancy-aided mixed convective flows past cylinders. Also, the point of inflexion found in the far-field flow is believed to be the reason for the unsteady plume at higher Ri.